OpenStudy (anonymous):

The lateral area of a cube is two hundred fifty-six square kilometers. What is its side length?

OpenStudy (anonymous):

Help

OpenStudy (raden):

let the length side of a square is x. the lateral area of a square = 4 * side * side known the lateral area of a square = 256 so, 4x^2 = 256 solve for x

OpenStudy (anonymous):

how

OpenStudy (raden):

4x^2 = 256 first, divide by 4 on both sides, what do you get ?

OpenStudy (anonymous):

64

OpenStudy (raden):

that shoul like this : 4x^2 = 256 4x^2 : 4 = 256 : 4 x^2 = 64

OpenStudy (raden):

now, take the suare root "right side" to get x

OpenStudy (raden):

* square root

OpenStudy (raden):